determine the integral of function ( x^3 -2 )(x^4 -8x)^0.5 dx



1
Expert's answer
2021-08-12T15:29:06-0400

Let us determine the integral


"\\int( \ufeff\ufeff\ufeffx^3 -2 )(x^4 -8x)^{0.5} dx\\\\\n=\\frac{1}{4}\\int( 4\ufeff\ufeff\ufeffx^3 -8 )(x^4 -8x)^{0.5} dx\\\\\n=\\frac{1}{4}\\int(x^4 -8x)^{0.5} d( x^4 -8x)\\\\\n=\\frac{1}{4}\\frac{2}{3}(x^4 -8x)^{\\frac{3}{2}}+C\\\\\n=\\frac{1}{6}(x^4 -8x)^{\\frac{3}{2}}+C"